OpenStudy (anonymous):

Find the x & y intercepts f(x)=x²-9x+7 please help me

OpenStudy (anonymous):

\[f(0)=0^2-9\times0+7=7\] which is the y intercept.\[f(x)=0=x^2-9x+7=>x=(9\pm \sqrt{53})/2\]Which are the x intercepts.

OpenStudy (anonymous):

the answer is x-intercept is (8.1,0.9) y-intercept is (0,7) how did they get these answers

OpenStudy (anonymous):

The x intercepts occur when y=0 (or, if you like, f(x)=0) so you just put this into you equation (remember, this is a quadratic function this will have two solutions and the function will therefore have two x intercepts). Similarly, the y intercept occurs when x=0, so you put this into your equation.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i know how to do y-intercepts but not x

OpenStudy (anonymous):

X intercepts occur when f(x)=0, so make your equation equal to zero and solve.

OpenStudy (anonymous):

plz show me how

OpenStudy (anonymous):

Do you know how to solve a quadratic equation?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

http://www.purplemath.com/modules/quadform.htm

OpenStudy (anonymous):

i just found a great website that did it for me ty

OpenStudy (anonymous):

Good :)